### v1

# Evaluating the impact of racial disproportionality in Stop & Search on expressive crimes in London

Abstract
Racial disproportionality in Stop & Search practices elicits substantial concerns about its societal and behavioral impacts. This paper aims to investigate the effect of disproportionality, particularly on the black community, on expressive crimes in London using data from January 2019 to December 2023. We focus on a semi-parametric partially linear structural regression method and introduce a Bayesian empirical likelihood procedure combined with double machine learning techniques to control for high-dimensional confounding and to accommodate the strong prior assumption. In addition, we show that the proposed procedure generates a valid posterior in terms of coverage. Applying this approach to the Stop & Search dataset, we find that racial disproportionality aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on expressive crimes.
Keywords: Stop & Search; Racial disproportionality; Partially linear regression; Bayesian estimation; Double machining methods
1 Introduction
1.1 Backgroud
Racial disproportionality in Stop & Search practices, particularly in cities such as London, raises important questions about its broader social and behavioral effects. Expressive crimes, such as vandalism, public disorder, or gang-related violence, are often argued to be influenced by social alienation, perceived injustice, or strained community-police relations, which can result from disproportionate policing practices.
Bowling and Phillips (2007) suggested that racial discrimination within policing is exacerbated by the Stop & Search powers in the UK targeting Black individuals disproportionately. The UK’s State of Policing report in 2022 highlighted that Stop & Search is an important tool for preventing and detecting crime, with significant public support when used fairly and proportionately, particularly in targeting weapons and drugs (HM Inspectorate of Constabulary and Fire & Rescue Services, 2022). Tiratelli et al. (2018) assessed the effectiveness of stop-and-search in reducing crime in London, with findings indicating some correlation between stop-and-search and drug offenses, but relatively limited impact on broader crime reduction.
In recent years, the issue of disproportionality in Stop & Search practices has remained a significant concern in London, where the disparities between ethnic groups have persisted despite various reform efforts. The March 2024 Disproportionality Board Data Pack (MOPAC, 2023), part of the Mayor’s Action Plan for Transparency, Accountability, and Trust in Policing, highlighted that Black individuals in London are still 3.3 times more likely to be stopped and searched compared to White individuals, and this figure rises to 6.2 times for searches related to weapons.
Scholars have identified that Stop & Search are highly concentrated in low-income neighborhoods or areas with higher populations of minority ethnic groups. According to Millner (2020), the increased use of new surveillance technologies, such as predictive policing software, has transformed policing in cities such as London, intensifying the monitoring of particular demographic groups. Suss and Oliveira (2023) and Meng (2017) demonstrated that the spatial patterns of Stop & Search in London are closely linked to areas characterized by higher levels of economic inequality and minority populations, revealing racial bias embedded in police practices. Oberwittler and Roché (2022) argued that in France, Germany and other European cities, police actions against adolescents in certain neighborhoods are often shaped by institutional biases tied to economic deprivation and race. Given these findings, it is crucial to use a data-driven approach to quantify the effect of racial disproportionality in Stop & Search. For example, a robust statistical approach is needed to account for the role of various confounding factors, such as differences in demographics across communities, socioeconomic conditions and policing priorities, from potential biases in police decision-making.
1.2 Statistical challenges
Based on the data from January 2019 to December 2023 in London aggregated within the borough level, the objective of this study is to estimate the effect of disproportionality in Stop & Search on expressive crimes. Estimation of the effect of disproportionality in Black people is complicated by the potential for confounding between the proportion of black population and the outcome. Confounding exists whenever exposure (or treatment) assignment is dependent on predictors that also influence the outcome, and appropriate adjustment is required to estimate the corresponding effect. Variables, such as the number of schools and proportion of the green space, can obscure the true relationship between these variables. For example, higher Stop & Search rates for Black people in areas might be attributed to an increased proportion of the black population in this borough, but it could also be influenced by other factors like the proportion of the educated household and the number of schools in the area. Determining this effect in this complex scenario requires careful consideration of variables and the implementation of rigorous research in data analysis. To estimate the effect of a policy, a common approach is to use partially linear regression as in Robinson (1988), which involves specifying separate regression models for both the treatment and outcome variables. The model for the treatment variable is often termed as ‘propensity score’ model. Propensity score adjustments (Rosenbaum and Rubin, 1983) have been extensively used to reduce confounding bias in estimating causal effects. The propensity score is defined as the conditional probability of the treatment assignment given confounding covariates. The propensity score can be used to break the dependence between confounders and exposure, to create balance in the distribution of confounders across exposure groups, and to facilitate correct inference. However, a key statistical challenge in applying partially linear regression for Stop & Search practices is the presence of high-dimensional confounders, where the parameter space for nuisance parameters grows with the sample size. In this case, the traditional semi-parameter theory might not offer valid inference for the parameter of interest. To overcome this challenge, Chernozhukov et al. (2018) proposed the use of machine learning methods to estimate the nuisance parameter and provided a simple and root-$n$ consistent procedure to estimate the parameter of interest via the Neyman moment equation and sample splitting.
Another challenge arises, when analyzing Stop & Search data, due to the strong prior beliefs and assumptions often held by policymakers and law enforcement agencies. These priors may reflect institutional perspectives on crime prevention strategies or operational practices, and we need to account for these prior beliefs in order to ensure robust and objective analysis, fostering evidence-based policy decisions. Therefore, a prior-to-posterior Bayesian analysis is preferable. The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the effect from a practical perspective. There is growing interest in the application of Bayesian methodology to two-stage propensity score regression analysis; however, most of existing Bayesian approaches require a full parametric specification for both structural equations in both regression models (see, for example McCandless et al., 2010; Kaplan and Chen, 2012). There are several works already attempting to solve this problem via a semi-parameter perspective (for example, Graham et al., 2016; Liu et al., 2020); these methods typically exploit the Bayesian bootstrap (Rubin, 1981; Chamberlain and Imbens, 2003) to perform inference. Recently, Luo et al. (2023) proposed to draw inference for the posterior from a Bayesian predictive distribution via a Dirichlet process model, extending the Bayesian bootstrap, and opening up the possibility to explore the new direction on doubly robust causal inference based on a non-parametric specification. In their work, the posterior samples are generated by resampling weights from the Pólya urn. It links with recent advancements in Bayesian empirical titled likelihood (Chib et al., 2018; Yiu et al., 2020; Luo et al., 2023), where the weights are replaced by the empirical probabilities. Antonelli et al. (2022) proposed to use the Gaussian process (GP) regression, and then the MCMC estimate is plugged into estimating equation. However, in present of high-dimensional confouders, those approaches can be slow when implementing Markov chain Monte Carlo (MCMC). In this paper, we aim to marry between the Bayesian method and the semi-parametric double machine learning method and provide a computationally efficient procedure to draw inference on the parameter of interest. We argue that our method bridges this gap by demonstrating how the approximate frequentist distribution theory can find an equally effective interpretation within a Bayesian framework in the high-dimensional setting. In addition, we show that the proposed procedure generates a valid posterior according to Monahan and Boos (1992), indicating a valid putative ‘posterior’ density computed by a non-standard method should still make probability statements consistent with Bayes’ rule.
1.3 Plan of paper
The remainder of this paper is organized as follows. In Section 2, we review the estimation procedure of partially linear regression, and introduce the notion of approximating Bayesian formulations and how to perform inference via the Neyman moment equation. We verify the validity of the proposed posterior inference in the spirit of Monahan and Boos (1992) in Section 2.3. Section 3 shows some simulation examples which compare the proposed method with some other frequentist and Bayesian approaches, following with Stop & Search data analysis in Section 4. Finally, Section 5 presents some concluding remarks and future research directions.
2 Statistical methodology
To estimate the effect of racial dispropotionality, we first consider the partially linear regression in Robinson (1988):
$Y=\mu(X)+\beta D+U,\;\;\;\mathbb{E}(U|X,D)=0,\;\;\;\mathbb{E}(D|X)=\pi(X)\$ (1)
where $Y$ is the outcome variable, $D$ is the treatment variable of interest (i.e., the proportion of Black people), and $X=\left(X_{1},\ldots,X_{p}\right)$ is a list of counfounders. Given independent and identically distributed data, $\left\{Z_{i}=(Y_{i},D_{i},X_{i}),i=1,\ldots,n\right\}$, the parameter of interest is $\beta$. If $D$ is exogenous conditional on $X$, then $\beta$ has the interpretation of the treatment effect parameter. Assuming there are no unmeasured confounders Under frequentist inference, the estimand of $\beta$ is consistent if $\mathbb{E}\left[Y\left|X,D\right.\right]$ is correctly specified. In addition, one can be constructed via specifying both outcome mean and propensity score models and use them in a combined estimation to yield a debiased estimator. To implement such inference, we need to specify the following two models:
  • Propensity score (PS) model: $\mathbb{E}(D|X)$, and it represents the conditional distribution of treatment given the confounders;
  • Outcome regression (OR) model: $\mathbb{E}\left(Y\left|X,D\right.\right)$, which represents the conditional distribution of the outcome given the treatment variable and confounder under the observational study.
As noted in Lee (2018), The model in (1) can be rewritten as a new regression model:
$Y-\mathbb{E}[Y|\pi(X)]=\beta[D-\pi(X)]+V,\;\;\;V=U+\mu(X)-\mathbb{E}[\mu(X)|\pi(X)]=0,\;\;\;V\perp\!\!\!\perp D|\pi(X)\$ (2)
and therefore $\mathbb{E}(V|\pi(X))=0$. The treatment effect parameter, $\beta$, can be estimated via the least square estimation procedure. This is equivalent to the Neyman-orthogonal moment equation noted in Chernozhukov et al. (2018), and $\beta$, is the solution $\mathbb{E}[\psi(Z;\beta)]=0$, where
$\psi(Z;\beta)=[D-\pi(X)][Y-\beta D-\mu(X)].$
In this way, we can reduce the high-dimensional problem to a single parameter problem if we assume $D$ is a scalar. This approach requires that we have access to both $\mu(X)$ and $\pi(X)$. The challenge arises in the high-dimensional setting, i.e., $p>n$; however, both of these functions can be estimated via machine learning (ML) methods. This double ML specification gives us flexibility to specify the structures of $\mu(X)$ and $\pi(X)$ while focusing on the treatment effect. However, it would be difficult to perform conventional Bayesian analysis with an strong prior input from the policy makers as there is in no distributional form for the conditional outcome mean and PS models. Conventional Bayesian inference focuses on updating prior belief in light of the data, and the data are summarized in the form of the likelihood. The relationship between prior beliefs and observable random quantities, $\mathbf{z}=\left(z_{1},\ldots,z_{n}\right)$, is formulated via the de Finetti representation, i.e.,
$f\left(z_{1},\ldots,z_{n}\right)=\int\prod_{i=1}^{n}f(z_{i}|\beta)\pi_{0}(\beta)d\beta.$
In the de Finetti representation, a full probabilistic model, $f(z_{i}|\beta)$, is required, and $\pi_{0}(\beta)$ is the prior belief about $\beta$. In this case, we have to examine procedures for Bayesian inference in the case where we wish to perform analysis of an approximate model, acknowledged to be misspecified compared to the data generating model. In the next section, we will seek solutions to incorporate the Neyman-orthogonal approach into a fully Bayesian procedure and demonstrate how the approximate frequentist distribution theory, which rests on moment constraint assumptions about distributions, can find an equally effective interpretation within a Bayesian framework.
2.1 Bayesian non-parametric procedure for the Neyman moment equation
[…]
If we specify $p$ as nonparametric, and use $\delta(f|p_{1},\ldots,p_{n})$ to measure the discrepancy between the true data generating model and the approximating model. Specifically, $p_{i}$ can be the solution of the dual formulation which satisfies certain moment conditions and can be summarized as the following constrained optimization problem
$\min_{p_{1},\ldots,p_{n}}\delta(f|p_{1},\ldots,p_{n})\text{ subject to }\sum_{i=1}^{n}p_{i}=1,\;\;p_{i}\geq 0,\;\;\sum_{i=1}^{n}p_{i}\psi(z_{i};\beta)=0.$ (4)
In light of Read and Cressie (2012), the the empirical Cressie-Read statistic can be used as a goodness-of-fit measure for discrete multivariate data. Then we specify $\delta(f|p_{1},\ldots,p_{n})\equiv\text{CR}(p)$, where
$\text{CR}(p)=\frac{2}{\lambda(1+\lambda)}\sum_{i=1}^{n}\left[(np_{i})^{-\lambda}-1\right],\;\;\;-\infty<\lambda<\infty$
where $\lambda$ is a user-specified parameter. From Baggerly (1998), this function can be rewritten as
$\text{CR}(p)=\left\{\begin{array}[]{cc}-2\sum_{i=1}^{n}\log(np_{i}),&\lambda=0\\ &\\ 2n\sum_{i=1}^{n}p_{i}\log(np_{i}),&\lambda=-1\\ &\\ \frac{2}{\lambda(1+\lambda)}\sum_{i=1}^{n}\left[(np_{i})^{-\lambda}-1\right],&\lambda\neq-1\text{ or }0.\\ \end{array}\right.$
Therefore, the solution in (4) is the nonparametric likelihood which seeks to reweight the sample so that it can also satisfy the moment condition (Qin and Lawless, 1994). It has been proven to possess many properties of the conventional parametric likelihood theory (Owen, 2001). A class of generalized empirical likelihood functions are studied in Imbens et al. (1998); Chernozhukov and Hong (2003); Newey and Smith (2004). In our example, we want to place a prior on the treatment effect parameter, $\beta$, directly, and update this prior in light of the data observed. Therefore, we can replace $f\left(z_{i}\left|\beta\right.\right)$ with the profile likelihood, $p_{i}$, and then the posterior distribution for $\beta$ becomes
$\pi\left(\beta\left|\mathbf{z}\right.\right)\propto\pi_{0}\left(\beta\right)\prod\limits_{i=1}^{n}p_{i}.$
In this way, we can incorporate a fully Bayesian procedure for $\beta$ while satisfying the conditions in (4). There are several choices of $\lambda$ leading to the existing empirical likelihood methods. For example, The case $\lambda=0$ yields the empirical likelihood (EL) case, where $p_{i}$ is obtained through the maximum likelihood estimation. When $\lambda=-1$, it yields the exponentially tilted empirical likelihood (ETEL) minimize, where $(p_{1},\ldots,p_{n})$ minimizes the Kullback-Leibler (KL) divergence between $(p_{1},\ldots,p_{n})$ and the empirical probabilities $(1/n,\ldots,1/n)$. This case has been extensively studied under model misspecification (Chib et al., 2018; Yiu et al., 2020; Luo et al., 2023). In addition, if $\lambda=-1/2$, it gives for the Hellinger distance (HD) measure discussed in Kitamura et al. (2013). As noted by Baggerly (1998), a unique solution exists for (4), provided that zero is inside the convex hull of $\psi(z_{i};\beta)$ for a given $\beta$. Applying the Lagrange multiplier approach, we can obtain the solution to (4) by minimizing
$\text{GEL}(p)=\frac{2}{\lambda(1+\lambda)}\sum_{i=1}^{n}\left[(np_{i})^{-\lambda}-1\right]+\kappa_{1}\left(\sum_{i=1}^{n}p_{i}-1\right)+n\kappa_{2}^{\top}\sum_{i=1}^{n}p_{i}\times\psi(z_{i};\beta)$
where $\kappa_{1}$ and $\kappa_{2}$ are the Lagrange multipliers. Setting $\partial\text{GEL}/\partial p_{i}=0$ yields the extreme of the form
$p_{i}=\left\{\begin{array}[]{cc}\frac{1}{n}\left[1+s+t\psi(z_{i};\beta)\right]^{-1/(1+\lambda)},&\lambda\neq-1\\ &\\ s\exp\left[t\psi(z_{i};\beta)\right],&\lambda=-1\\ \end{array}\right.$ (5)
where $s$ and $t$ are normalized constant and determined by
$\frac{1}{n}\sum_{i=1}^{n}\left(1+s+t\psi(z_{i};\beta)\right)^{-1/(1+\lambda)}=1,$
$\frac{1}{n}\sum_{i=1}^{n}\left(1+s+t\psi(z_{i};\beta)\right)^{-1/(1+\lambda)}\psi(z_{i};\beta)=0.$
Therefore, by substituting (5) into the posterior distribution, we obtain the following posterior distribution
$\pi(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\left\{\begin{array}[]{cc}\prod_{i=1}^{n}\frac{1}{n}\left[1+\hat{s}(z,\beta)+\hat{t}(z,\beta)\psi(z_{i};\beta)\right]^{-1/(1+\lambda)}&\lambda\neq-1\\ &\\ \prod_{i=1}^{n}\frac{\exp\left(\hat{t}(z,\beta)^{\top}\psi(z_{i};\beta)\right)}{\sum_{j=1}^{n}\exp\left(\hat{t}(z,\beta)\psi(z_{j};\beta)\right)}&\lambda=-1\\ \end{array}.\right.$ (6)
The following algorithm summarizes the computation step to the Bayesian generalized empirical likelihood method.
Algorithm 1 Algorithm to obtain the posterior sample of $\beta$ via the Bayesian generalized empirical likelihood.
1:$\mathcal{D}=(z_{1},\ldots,z_{n})$
2:Estimate $\pi(x)$ and $\mu(x)$ using some ML methods.
3:for $j$ to $1:J$ do
4:     Sample $\beta^{(j)}\sim\pi_{j}(\beta\left|\mathbf{z}\right.)$ using the MCMC approach, where $\pi_{j}(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\prod_{i=1}^{n}p_{i}^{(j)}$.
5:     $\left(p^{(j)}_{1},\ldots,p^{(j)}_{n}\right)$ is the solution to $\min_{p_{1},\ldots,p_{n}}\text{CR}(p)\text{ subject to }\sum_{i=1}^{n}p_{i}=1,\;\;p_{i}\geq 0,\;\;\sum_{i=1}^{n}p_{i}\psi(z_{i};\beta^{(j-1)})=0$
6:end for
7:return $\left(\beta^{(1)},\ldots,\beta^{(J)}\right)$.
2.2 The role of sample splitting procedures
In Chernozhukov et al. (2018), they found that sample splitting plays a key role in reducing the bias when estimating $\beta$. The remainder term in the Taylor expansion involves the product between the error term in the PS model and the bias using ML method to estimate $\mu(\cdot)$. In some cases, this remainder term might not vanish when $n\to\infty$ as the two terms are correlated. In conventional semi-parametric analysis, we can impose Donsker conditions to restrict the class of functions that contains the estimator of $\mu(\cdot)$ so that the remainder term will be negligible. But when using ML methods where $p$ is modelled as $n$ increases, Donsker conditions are inappropriate. In Figure 1, we replicate the example in Chernozhukov et al. (2018), with $\hat{\mu}(X)=\mu(X)+(Y-\mu(X))/{n^{1/3}}$ and $\hat{\pi}(X)=\pi(X)$. We use the full sample to estimate $\beta$ using (1) and Algorithm 1 with a vague prior $\mathcal{N}(0,10000)$ and $\lambda=0,-1,-1/2$, which corresponds to EL, ETEL and HD cases respectively. The first column of Figure 1 shows the results of 10,000 replicates, and the histograms of the posterior mean via the Bayesian method in Algorithm 1. In this simple example, both methods indicate some biases from the full sample approach while the Bayesian method has a slightly smaller bias. In light of this consideration, Chernozhukov et al. (2018) proposed the use of sample splitting, that is, the data are partitioned into $K$ groups. The functions $\hat{\mu}_{k}(\cdot)$ and $\hat{\pi}_{k}(\cdot)$ are estimated using all the data excluding the $k$th group. Then the double ML estimator for $\beta$ is the solution to $01/K\sum_{k=1}^{K}\mathbb{E}_{k}[\psi(Z;\beta)]=$ 0, where $\mathbb{E}_{k}(\cdot)$ is the empirical expectation over the $k$th fold of the data. This creates the independence between two terms, leading to unbiased estimation. Therefore, it is necessary to amend Algorithm 1 to incorporate the sample splitting strategy to remove the bias. In essence, we can mimic the procedure to partition the data into $K$ groups and estimate the $\mu_{k}(\cdot)$ and $\pi_{k}(\cdot)$ using all the data excluding the $k$th group and then use them to obtain the non-parametric probability $p_{i}$ in (4) with $i\in$ $k$th group only. Algorithm 2 summarizes the update algorithm to generate the posterior sample using sample splitting.
Algorithm 2 Algorithm to obtain the posterior sample of $\beta$ via the Bayesian generalized empirical likelihood with sampling splitting.
1:$\mathcal{D}=(z_{1},\ldots,z_{n})$
2:Partition the data into $K$ groups (roughly equal size), $\mathcal{D}=\left\{\mathcal{D}_{1},\ldots,\mathcal{D}_{K}\right\}$.
3:for $j$ to $1:J$ do
4:     for $k$ to $1:K$ do
5:         Estimate $\mu_{k}(\cdot)$ and $\pi_{k}(\cdot)$ using some ML methods with data $\mathcal{D}\setminus\mathcal{D}_{k}$.
6:         Obtain the non-parametric probability, $\left\{p^{(j)}_{i}\right\}_{i\in\mathcal{D}_{k}}$, by solving the optimization problem in (4) with $\psi(z_{i};\beta^{(j-1)})$ for $i\in\mathcal{D}_{k}$.
7:     end for
8:     Sample $\beta^{(j)}\sim\pi_{j}(\beta\left|\mathbf{z}\right.)$ using the MCMC approach, where $\pi_{j}(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\prod_{i=1}^{n}p_{i}^{(j)}$.
9:end for
10:return $\left(\beta^{(1)},\ldots,\beta^{(J)}\right)$.
The second column of Figure 1 displays the results using two-fold sample splitting, i.e., $K=2$. Both the double machine learning (ML) methods and the amended Bayesian approach effectively eliminate bias from the full-sample estimation procedure. The posterior distribution obtained using Algorithm 2 exhibits similar distributional properties across 10,000 replications.
Figure 1: Comparison between double ML (DML) and Bayesian empirical likelihood methods using full-sample and sample splitting approaches ($n=500$). The red curve represents the standard normal density.
2.3 Validating the posterior uncertainty
In this section, we evaluate whether the posterior in (6) is a valid posterior inference in terms of uncertainty. As demonstrated in Chernozhukov et al. (2018), if the estimated functions of $\mu(x)$ and $\pi(x)$ converge to the true functions in probability under some regularity conditions, the estimator for $\beta$ in (2) will converge in distribution to a normal distribution, centered at $\beta$ with variance, $\mathbb{E}[(D-\pi(X))^{2}]^{-1}\mathbb{E}[(Y-\beta D-\mu(X))^{2}]$. This is the semi-parametric efficiency bound for $\beta$.
The ‘posterior’ density in Algorithm 2, $\pi(\beta|z_{1:n})$, is a non-standard posterior inference as the likelihood is replaced by the profile likelihood with plug-in estimates for $\mu(x)$ and $\pi(x)$. Despite the theoretical support of the ETEL case in Schennach (2007); Chib et al. (2018), one would want to assess if this plug-in coupled with sample-splitting strategy is valid Bayesian inference. That is, if interval $S_{\alpha}(z)$ is a designated $1-\alpha$ probability interval, a ‘posterior’ density in Algorithm 2, $\pi(\beta|z_{1:n})$, will have the property $\mathbb{P}_{\pi}\left[\beta\in S_{\alpha}(z))\right]=1-\alpha$, if $Z_{1},\ldots,Z_{n}$ are drawn from the true data generating model. Monahan and Boos (1992) proposed a notion of proper Bayesian inference by replacing the parametric likelihood with an alternative likelihood function. They stated that the ‘posterior’ density, which derives from the alternative likelihood, should follow the law of the probability deriving from the Bayes’s rule. The posterior density is defined as valid by coverage if $\mathbb{P}_{\pi}\left[\beta\in S_{\alpha}(z))\right]=1-\alpha$, if $Z_{1},\ldots,Z_{n}$ are drawn from the true data generating model. The posterior coverage set, $S_{\alpha}(z)$, resulting from a valid posterior, should achieve nominal coverage under the joint measure of $Z$ and $\theta$, that is, $P_{\pi}(\beta\in S_{\alpha}(z))$ should have expectation $1-\alpha$ for data generated under the measure $\pi_{0}(\beta)f(z|\beta)$ on $(\beta,Z)$ for every absolutely continuous prior, $\pi_{0}(\ldotp)$. To verify this property, let
[…]
and if $\pi(\beta|\mathbf{z})$ is a valid posterior, then $H$ follows Uniform$(0,1)$. In practice, if we generate $\beta_{k}(k=1,\ldots,m)\sim\pi_{0}(\ldotp)$ and the data, $z_{1:n}^{(k)}$, from $f(\ldotp|\beta_{k})$, and compute the posterior according to Algorithm 2. Then we can obtain $H_{k}$ based on (7) by replacing $\beta$ with $\beta_{k}$. If the distribution of $H_{k}$ follows the uniform distribution, then the posterior distribution generated from Algorithm 2 is defined as a coverage proper posterior, yielding valid posterior inference. This evaluation method for the validity of posterior inference has also been applied to verify the correctness of Bayesian computation (see for example, Talts et al., 2018).
In light of this approach, we investigate the validity of the proposed generalized empirical likelihood approach via a simulation study using the same set-up with the sample-splitting simulation, with each $\beta_{k}$ ($k=1,\ldots,10000)$ generating from $N(1,2)$. Figure 2 contains the histograms of the simulated $H$ over 10000 simulation runs with the associated $p$-values of the Kolmogorov–Smirnov test for uniformity. For all cases, it suggests that simulation $H$ values follow uniform distribution and indicates proper posterior inference according to Monahan and Boos (1992) when $\lambda=0,-1,-1/2$. Therefore, this simulation result gives us evidence that the proposed method according to Algorithm 2 can be regarded as a valid approach for posterior inference in terms of coverage.
Figure 2: Histgrams for $H$ statistics using the Bayesian generalized empirical likelihood. $P$-values represents the Kolmogorov–Smirnov test for uniformity of $H$.
3 Simulation
In this section, we examine the performance of the proposed Bayesian method described in the previous section. We consider the following models for our simulation studies.
  • EL, ETEL, HD: The Bayesian generalized empirical likelihood with $\lambda=0,-1,-1/2$ respectively described in Section 2.1, with calculation using Alogrithm 2.
  • BDR-HD: The Bayesian doubly robust high-dimension method proposed in Antonelli et al. (2022), where the propensity score and outcome are estimated via regression models with the Gaussian process prior, and then the MCMC estimate is plugged in to a doubly robust estimator. The variance is adjusted through the frequentist bootstrap so that it will achieve the nominal coverage rate.
  • DML: The frequentist double machine learning approach proposed in Chernozhukov et al. (2018).
For all the Bayesian methods, we generate $5,000$ MCMC samples and $1,000$ burn-in iterations each simulation with $1,000$ simulation replicates.
3.1 Binary exposure
In this case, we consider $D$ is binary and simulate $X=(X_{1},X_{2},\ldots,X_{p})\sim\mathcal{N}_{p}\left(0,\Sigma\right)$ and $\Sigma_{ij}=1$ if $i=j$ and 0.3 otherwise, and then simulate
[…]
$Y\left|D,X\right.$ $\sim\mathcal{N}\left(D+0.5X_{1}+X_{3}-0.1X_{4}-0.2X_{7},1\right).$
We are interested in estimating the average treatment effect (ATE), i.e, $\beta$. In the analyses, we take $p=500$ and $n=50,200$ and place a non-informative prior, $\mathcal{N}(0,10000)$, for the ATE. Table 1 shows the performance of ETEL, EL, HD, DML across machine learning approaches (Lasso, Random Forest, and Neural Network) and BDR-HD in terms of bias, root mean squared error (RMSE), and coverage rate. For the proposed Bayesian method, it maintain low bias and RMSE with relatively high coverage rates. Specifically, EL and HD generally achieve slightly better precision than ETEL, with HD sometimes offering the lowest bias. In terms of machine learning approaches, neural networks, although effective in reducing bias, show higher RMSE and lower coverage rates compared to Lasso and Random Forest. DML method demonstrates quite similar results with the proposed Bayesian method, in line with our previous demonstration about the uncertainty quantification. BDR-HD shows similar bias as other methods, but it achieves the lowest RMSE and the highest coverage rate due to the extra bootstrap step to adjust the posterior variance. However, the running time per iteration for $n=50$ of BDR-HD is approximately five times longer than that of the proposed Bayesian method. Overall, the proposed method demonstrates a reliable balance between the statistical performances and computational intensity.
Table 1: Binary exposure: Simulation results of the marginal causal effect under high-dimensional settings, with true value equal to 1, on 1000 simulation runs on generated datasets of size $n$.
$n=50$ $n=200$
Bias RMSE Coverage rate (%) Bias RMSE Coverage rate (%)
ETEL (Lasso) 0.08 0.42 93.1 0.08 0.19 93.4
EL (Lasso) 0.07 0.42 91.7 0.07 0.19 92.1
HD (Lasso) 0.08 0.42 92.8 0.04 0.18 91.8
ETEL (Random forest) 0.06 0.43 90.9 0.08 0.19 92.4
EL (Random forest) 0.06 0.42 92.8 0.07 0.20 92.0
HD (Random forest) 0.07 0.43 90.9 0.08 0.19 92.7
ETEL (Neural network) 0.08 0.47 87.6 0.03 0.22 87.4
EL (Neural network) 0.06 0.44 90.5 0.04 0.23 88.3
HD (Neural network) 0.01 0.46 90.7 0.02 0.22 90.0
DML (Lasso) 0.10 0.41 92.7 0.05 0.17 93.1
DML (Random forest) 0.09 0.41 93.9 0.08 0.20 92.0
DML (Neural network) 0.08 0.43 90.2 0.05 0.22 88.8
BDR-HD 0.09 0.36 97.7 0.07 0.15 95.4
3.2 Continuous exposure
In this example, we consider $D$ is continuous and simulate $X=(X_{1},X_{2},\ldots,X_{p})\sim\mathcal{N}_{p}\left(0,\Sigma\right)$ and $\Sigma_{ij}=1$ if $i=j$ and 0.05 otherwise, and then simulate
$D\left|X\right.$ $\sim\mathcal{N}\left(0.45X_{1}+0.9X_{2}-0.4X_{5}\right)$
$Y\left|D,X\right.$ $\sim\mathcal{N}\left(D+0.5X_{1}+X_{3}-0.1X_{4}-0.2X_{7},1\right).$
In this example, the coefficient associated with $D$ is the main regression coefficient and is the parameter of interest. We set $p=40$ and $n=40$, and in the proposed Bayesian method, we assign an informative prior, $\mathcal{N}(1,2)$, to $\beta$. Table 2 shows the results of bias, RMSE and coverage rate across all methods. Most of the methods perform well in terms of bias, ranging between $-0.01$ and $0.06$, while the RMSE varies slightly, with most values in between $0.13$ and $0.19$. In particular, Lasso-based approaches have very low bias ($0.01-0.02$), while random forest based methods show slightly higher bias. In terms of RMSE, the proposed Bayesian method displays smallest RMSEs $(0.13-0.14)$ across all ML methods. DML methods showing a slightly higher RMSE $(0.16-0.19)$, while BDR-HD has a much higher RMSE. Regarding coverage rates, EL-based methods always achieve higher rates $(90.6-94.0)$, while ETEL-based methods show a bit lower coverage $(85.8-89.9)$. BDR-HD reaches the coverage rate at 95.9, closest to the nominal level. We also notice that methods using neural network generally have lower coverage rates than methods using Lasso and random forest. Results indicate that the proposed Bayesian method, coupled with an informative prior, performs well overall in terms of balancing low bias, low RMSE, and adequate coverage rates.
Table 2: Continuous exposure: Simulation results of the marginal causal effect under high-dimensional settings, with true value equal to 1, on 1000 simulation runs with $n=40$ and $p=40$.
Bias RMSE Coverage rate (%)
ETEL (Lasso) 0.02 0.13 89.6
EL (Lasso) 0.02 0.13 94.0
HD (Lasso) 0.01 0.13 91.7
ETEL (Random forest) 0.06 0.13 89.9
EL (Random forest) 0.06 0.13 93.0
HD (Random forest) 0.06 0.13 91.5
ETEL (Neural network) 0.04 0.14 85.8
EL (Neural network) 0.04 0.14 90.6
HD (Neural network) 0.04 0.14 88.0
DML (Lasso) -0.01 0.19 90.2
DML (Random forest) 0.02 0.16 92.3
DML (Neural network) 0.01 0.18 88.5
BDR-HD 0.03 0.75 95.9
4 Application: Stop and Search in London, UK
4.1 Stop and Search data
In this section, we apply the proposed methodology to the Stop & Search data in London to enhance our understanding of the impact of the racial disproportionality, particularly on expressive crimes. The Stop & Search dataset consists of individual stop and search monthly records in London from January 2019 to December 2023, including date and time, street-level location, ethnicity, gender and age of the person stopped, legislation, object searched and outcome. The data have been sourced from the Metropolitan Police and City of London police force and published in the Single Online Home National Digital Team under Open Government Licence v3.0. They are available openly from https://data.police.uk/docs/method/stops-at-location/. Before publishing, the location coordinates of the stop are anonymized (detailed methods of anonymization from the publisher can be found at https://data.police.uk/about/#location-anonymisation and the age of the person stopped has been adjusted to a corresponding age group (e.g. 18-24).
Category Searched Objects Legislation
Acquisitive objects ‘Stolen goods’, ‘Article for use in theft’ ‘Police and Criminal Evidence Act 1984 (section 1)’, ‘Criminal Justice Act 1988 (section 139B)’, ‘Police and Criminal Evidence Act 1984 (section 6)’
Expressive objects ‘Offensive weapons’, ‘Anything to threaten or harm anyone’, ‘Firearms’, ‘Fireworks’, ‘Evidence of offences under the Act’, ‘Crossbows’, ‘Game or poaching equipment’ ‘Police and Criminal Evidence Act 1984 (section 1)’, ’Criminal Justice Act 1988 (section 139B)’, ‘Police and Criminal Evidence Act 1984 (section 6)’, ‘Criminal Justice and Public Order Act 1994 (section 60)’, ‘Firearms Act 1968 (section 47)’
Drug objects ‘Controlled drugs’, ‘Psychoactive substances’ ‘Misuse of Drugs Act 1971 (section 23)’, ‘Psychoactive Substances Act 2016 (s36(2))’
In this research, 2.6 million Stop & Search incidents have been compiled for the observation period from January 2019 to December 2023, on seven legislative grounds (Legislation in the above table) for 11 types of Searched Objects. Upon the aggregation of the data into three listed categories, drug objects related stop and search incidents took up 64% over the observation period, as compared to 12% for acquisitive objects and 24% for expressive objects. The searched subjects’ ethnic information have been collected from “officer defined ethnic group”.
4.2 Demographic data on ethnic groups
The demographic data are taken from Census 2021 data by the Office for National Statistics (ONS) and are available from https://www.nomisweb.co.uk/sources/census_2021. The demographic data features have been collocated at local authority, “borough”, level as the analytical geographical unit in this research.
The ethnic groups in Census 2021 are mainly 5 groups including “Asian, Asian British, Asian Welsh”, “Black, Black British, Black Welsh, Caribbean or African”, “Mixed or Multiple”, “White”, and ”Other ethnic group” (ONS, 2021). To be consistent with the categories of ethnic groups in the Stop & Search dataset, we further group the ethnic groups into four groups as: “Asian, Asian British, Asian Welsh”, “Black, Black British, Black Welsh, Caribbean or African”, “White”, and “Others” (including Mixed and Other groups in the official Census dataset). From the latest Census 2021 data, ethnic groups’ proportions in London are unevenly distributed in that, White people take up 53.73%, followed by Asian people at 20.7%, Black people at 13.52%, and Others at 12.05%.
Exploratory data analysis on Stop and Search subjects’ demographics, especially the ethnic compositions, has found disproportionally distributed Stop & Search among the ethnic groups. White people took up on average 40% over time (ranged from 38% to 42%), Black people took up on average 38% during the observation period (ranged from 36% to 40%), Asian people took up on average 17% (ranged from 14% to 18%), then Others at 5% (ranged from 4% to 6%), as depicted in Figure 1.
Figure 3: Stop and Search Subjects Ethnic Groups Proportions
4.3 Outcome definition
To apply our method, we define the outcome variable as the dispropotionality index (DI) for expression crime targeting Black people in each Borough (total 33 Boroughs) and it can be calculated as:
$\text{DI}_{i}=\frac{\text{rate of black people in Stop \& Search for expression crime in Borough }i}{\text{rate of average black people in Stop \& Search for expression crime in London}}.$
This index compares the rate of Black people in the Stop & Search for expression crime in each Borough to the average rate in London:
  • If $\text{DI}=1$, it indicates no disproportionality;
  • If $\text{DI}<1$, it indicates too little Stop & Search below expectation;
  • If $\text{DI}>1$, it indicates too much Stop & Search above expectation.
Figure 4: DI for Black People in Stop and Search for Expressive Reasons
Boroughs in South-West London, such as Richmond upon Thames and Kingston upon Thames, as well as City of London, are depicted in dark red with highest DI; whilst boroughs in the east alongside the banks, such as Lewisham, Greenwich, Bexley, Newham, Barking and Dagenham, as well as the very North Enfield and far South Croydon, have presented lower DI among boroughs.
Figure 5: Monthly Stop and Search in London (2019-2023)
4.4 Results
We use the percentage of black people in each borough as the treatment variable. The estimated effect is adjusted by variables such as percentage of males (gender), unemployment number and migrant rate, with a total of 31 variables. A summary of all variables, including their descriptive statistics aggregated in the borough level in the analysis, is given in Table 3. We consider the Bayesian partially linear structural equation models with an informative prior, $\mathcal{N}(0,1)$, indicating the prior belief and assumption on no racial disproportion in Stop & Search practice in London. We implement the Bayesian generalized empirical likelihood with $\lambda=0,-1,-1/2$ coupled with Lasso, random forest and neural network methods, and generate in total 10000 posterior samples in each case with 1000 burn-in iterations. Table 4 shows the results from applying the proposed method. Across all cases, the posterior mean estimates are consistently negative, ranging from -0.55 to -0.31. In addition, while it shows negligible differences across ETEL, EL and HD, the results for different machine learning methods vary, especially the width of the credible intervals. Those for Lasso are the widest, and this might be due to the high correlation among some features, leading to high variability in cross-validation when selecting the optimal tuning parameter. Random forest and neural network models have relatively narrower confidence intervals, with the narrowest under EL and HD methods among them. Overall, all methods suggest a negative relationship in terms of posterior mean, indicating that disproportionality in Stop & Search aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on Stop & Search against Black people for expressive crimes. That is, the higher the percentage of black residents composition in the borough, the lower the predicted value of DI of Stop & Search on Black people for expressive reasons.
Table 3: Descriptive statistics of the Stop & Search data in London
Variable Min. Max. Mean Std. Dev.
Dispropotionality index 1.66 12.58 4.12 2.46
Proportion of Black population 1.89 17.59 12.64 6.92
Population Density 2198 15703 7291 3670.89
Proportion of Male Residents 0.47 0.55 0.49 0.01
Proportion of Migrant Residents 0.57 6.42 1.94 1.30
Number of unemployment 2653 123179 83149 24860.92
Proportion of Residents Self-claimed as Unhealthy 2.72 5.55 4.23 0.63
Proportion of Disabled Residents 21.41 32.38 26.41 2.37
Proportion of Students Residents 13.90 28.54 21.90 2.53
Proportion of Households without Cars 21.53 77.20 42.90 16.71
Proportion of Households in Renting 29.54 74.26 53.37 13.84
Household Room Occupation Rate 9.39 45.02 27.40 7.62
Proportion of Residents with Higher Education Degree 29.52 74.18 47.95 9.93
Proportion of Deprived Households 38.96 62.41 51.46 5.45
Proportion of Residents Lacking Care Supports 91.30 93.70 92.29 0.62
Proportion of Households Lack of Family Cohension 33.01 54.54 42.87 5.61
Proportion of Households Living in Unstable Status 13.93 43.27 30.55 6.34
Proportion of Greenspace Area 0.01 0.487 0.17 0.12
Proportion of Young Residents Under 18 0.28 0.49 0.37 0.05
Areas (km2) 2.90 150.14 47.68 32.75
Density of Roads 61.23 258.62 128.08 41.78
Density of Manufacturing Stores and Places 3.00 94.00 13.82 16.15
Number of Residential Places 420 3816 1094 603.35
Density of Residential Places 6.00 401.00 47.88 73.89
Number of Manufacturing Stores and Places 195 730 401 127.83
Number of Public Transport Stations 282.00 1821.00 957.40 353.16
Density of Public Transport Stations 10.00 97.00 13.82 16.02
Number of Pubs 28.00 447.00 123.10 82.46
Density of Pubs 0.62 76.10 6.37 13.39
Number of Retail Stores 486 3536 1282 522.86
Density of Retail Stores 8.00 167.00 44.94 42.29
Number of Schools 8.00 171.00 102.20 29.56
Density of Schools 0.85 8.00 2.98 1.83
Table 4: Posterior mean of the effect of for expression crime targeting black population with associated 95% credible interval.
Lasso Random forest Neural network
ETEL $-0.59$ $-0.31$ $-0.41$
$(-2.17,0.95)$ $(-0.52,-0.14)$ $(-0.73,-0.16)$
EL $-0.56$ $-0.31$ $-0.40$
$(-2.21,0.99)$ $(-0.53,-0.14)$ $(-0.78,-0.14)$
HD $-0.55$ $-0.31$ $-0.41$
$(-2.18,1.00)$ $(-0.55,-0.14)$ $(-0.78,-0.16)$
5 Discussion
In this paper, we introduce a Bayesian procedure for semi-parametric partially linear structural regression which allows us to place a specific prior to the parameter of interest. In addition, this approach also addresses issues in high-dimensional scenarios. By integrating double machine learning method and sample splitting procedure into our Bayesian framework, it preserves key asymptotic properties and consistent estimation on the parameter of interest. In particular, we showed that computations following this paradigm yield valid posterior inference in terms of coverage according to Monahan and Boos (1992). Our approach also accommodates flexible machine learning models for treatment and outcome, mitigating the impact of model misspecification. In our application, we estimate the effect of impact of ethnic disproportionality in Stop & Search on expressive crimes in London with an informative prior, and conclude that: Stop & Search practices in London are disproportionally distributed among London boroughs and disproportionality targeting the Black community is mitigated by considering the proportion of the Black population for expressive crimes. The ethnic composition of demographics plays a significant role in affecting the identified disproportionality, with lower DI alleviated by higher percentage of black population in target boroughs.
Bayesian methods hold significant interest in applied scientific causal research. These methods enable direct probability statements regarding treatment effectiveness and facilitate sensitivity assessments with different prior or expert inputs. There remains many scopes for future research to explore various possibilities in Bayesian semi-parametric inference in high-dimensional settings. For example, the proposed methodology can be widely applied in other causal settings when the traditional Bayesian set-up requires over-specifying the model condition. Moreover, this Bayesian method, coupled with machine learning methods, also presents an important direction for advancement to accommodate complex data structures, such as time-varying treatments, longitudinal data, or hierarchical frameworks in high-dimensional scenarios.
Acknowledgments
This research has been supported by the Engineering & Physical Sciences Research Council (EP/Y029755/1). The authors would like to thank the Evidence and Insight team from MOPAC (Mayor’s Office for Policing and Crime, Greater London Authority) for their valuable comments and suggestions that contributed to the progress of this research.

### v2

@@
-Evaluating the impact of racial disproportionality in Stop & Search on expressive crimes in London
+A scalable Bayesian double machine learning framework for high dimensional causal estimation, with application to racial disproportionality assessment
 Abstract
-Racial disproportionality in Stop & Search practices elicits substantial concerns about its societal and behavioral impacts.
+Racial disproportionality in Stop and Search practices elicits substantial concerns about its societal and behavioral impacts.
 This paper aims to investigate the effect of disproportionality, particularly on the black community, on expressive crimes in London using data from January 2019 to December 2023.
-We focus on a semi-parametric partially linear structural regression method and introduce a Bayesian empirical likelihood procedure combined with double machine learning techniques to control for high-dimensional confounding and to accommodate the strong prior assumption.
-In addition, we show that the proposed procedure generates a valid posterior in terms of coverage.
-Applying this approach to the Stop & Search dataset, we find that racial disproportionality aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on expressive crimes.
-Keywords: Stop & Search; Racial disproportionality; Partially linear regression; Bayesian estimation; Double machining methods
+We focus on a semi-parametric partially linear structural regression method and introduce a scalable Bayesian empirical likelihood procedure combined with double machine learning techniques to control for high-dimensional confounding and to accommodate strong prior assumptions.
+In addition, we show that the proposed procedure yields a valid posterior in terms of coverage.
+Applying this approach to the Stop and Search dataset, we find that racial disproportionality aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on expressive crimes.
+Keywords: Partially linear regression; Bayesian estimation; Double machining methods; Stop and search; Racial disproportionality
 1 Introduction
-1.1 Backgroud
-Racial disproportionality in Stop & Search practices, particularly in cities such as London, raises important questions about its broader social and behavioral effects.
+1.1 Background
+Racial disproportionality in stop and search practices, particularly in cities such as London, raises important questions about its broader social and behavioral effects.
 Expressive crimes, such as vandalism, public disorder, or gang-related violence, are often argued to be influenced by social alienation, perceived injustice, or strained community-police relations, which can result from disproportionate policing practices.
-Bowling and Phillips (2007) suggested that racial discrimination within policing is exacerbated by the Stop & Search powers in the UK targeting Black individuals disproportionately.
-The UK’s State of Policing report in 2022 highlighted that Stop & Search is an important tool for preventing and detecting crime, with significant public support when used fairly and proportionately, particularly in targeting weapons and drugs (HM Inspectorate of Constabulary and Fire & Rescue Services, 2022).
+Bowling and Phillips (2007) suggested that racial discrimination within policing is exacerbated by the stop and search police powers in the UK targeting Black individuals disproportionately.
+The UK’s State of Policing report in 2022 highlighted that stop and search is an important tool for preventing and detecting crime, with significant public support when used fairly and proportionately, particularly in targeting weapons and drugs (HM Inspectorate of Constabulary and Fire & Rescue Services, 2022).
 Tiratelli et al.
-(2018) assessed the effectiveness of stop-and-search in reducing crime in London, with findings indicating some correlation between stop-and-search and drug offenses, but relatively limited impact on broader crime reduction.
-In recent years, the issue of disproportionality in Stop & Search practices has remained a significant concern in London, where the disparities between ethnic groups have persisted despite various reform efforts.
+(2018) assessed the effectiveness of stop and search in reducing crime in London, with findings indicating some correlation between stop and search and drug offenses, but relatively limited impact on broader crime reduction.
+In recent years, the issue of disproportionality in stop and search practices has remained a significant concern in London, where the disparities between ethnic groups have persisted despite various reform efforts.
 The March 2024 Disproportionality Board Data Pack (MOPAC, 2023), part of the Mayor’s Action Plan for Transparency, Accountability, and Trust in Policing, highlighted that Black individuals in London are still 3.3 times more likely to be stopped and searched compared to White individuals, and this figure rises to 6.2 times for searches related to weapons.
-Scholars have identified that Stop & Search are highly concentrated in low-income neighborhoods or areas with higher populations of minority ethnic groups.
+Scholars have identified that stop and search events are highly concentrated in low-income neighborhoods or areas with higher populations of minority ethnic groups.
 According to Millner (2020), the increased use of new surveillance technologies, such as predictive policing software, has transformed policing in cities such as London, intensifying the monitoring of particular demographic groups.
-Suss and Oliveira (2023) and Meng (2017) demonstrated that the spatial patterns of Stop & Search in London are closely linked to areas characterized by higher levels of economic inequality and minority populations, revealing racial bias embedded in police practices.
+Suss and Oliveira (2023) and Meng (2017) demonstrated that the spatial patterns of stop and search in London are closely linked to areas characterized by higher levels of economic inequality and minority populations, revealing racial bias embedded in police practices.
 Oberwittler and Roché (2022) argued that in France, Germany and other European cities, police actions against adolescents in certain neighborhoods are often shaped by institutional biases tied to economic deprivation and race.
@@
 For example, a robust statistical approach is needed to account for the role of various confounding factors, such as differences in demographics across communities, socioeconomic conditions and policing priorities, from potential biases in police decision-making.
-1.2 Statistical challenges
-Based on the data from January 2019 to December 2023 in London aggregated within the borough level, the objective of this study is to estimate the effect of disproportionality in Stop & Search on expressive crimes.
-Estimation of the effect of disproportionality in Black people is complicated by the potential for confounding between the proportion of black population and the outcome.
+1.2 Statistical challenges and related work
+Disparities in institutional decision-making have significant implications for both individual well-being and broader societal outcomes.
+In healthcare, unequal access to treatments or disproportionality in clinical decision-making across population groups can lead to persistent health inequities.
+Similarly, differential treatment by law enforcement in policing may raises important questions about its broader social and behavioral effects.
+Quantifying racial bias in policing has become a central question in policy-oriented social science and applied statistics.
+Despite wide availability of administrative policing data, such as stop records, arrest databases, and officer-reported activity logs, using these data to make credible causal inferences about discrimination remains methodologically challenging.
+Knox et al.
+(2020) demonstrate that administrative policing datasets inherently condition on the outcome of police discretion, and show that estimands targeting racial bias that condition on being stopped can potentially have the opposite signs as the estimands of interest.
+A policing agency could appear unbiased or even favorable to minority groups among stopped individuals while simultaneously discriminating in its choice of whom to stop.
+Based on this, Zhao et al.
+(2022) clarify the definition of discrimination in causal inference terms and emphasize the need for estimands that avoid conditioning on post-treatment variables.
+Their work stresses that identifying racial disparities requires not only conditioning strategies but careful articulation of what causal question is being answered.
+From a causal inference perspective, estimation of racial disproportionality is complicated by confounding of the exposure-outcome association.
 Confounding exists whenever exposure (or treatment) assignment is dependent on predictors that also influence the outcome, and appropriate adjustment is required to estimate the corresponding effect.
-Variables, such as the number of schools and proportion of the green space, can obscure the true relationship between these variables.
-For example, higher Stop & Search rates for Black people in areas might be attributed to an increased proportion of the black population in this borough, but it could also be influenced by other factors like the proportion of the educated household and the number of schools in the area.
+Variables, such as the number of schools and proportion of green space, can obscure the true relationship between these variables.
+For example, higher stop and search rates for Black people in areas might be attributed to an increased proportion of the black population in this borough, but it could also be influenced by other factors like the proportion of the educated household and the number of schools in the area.
 Determining this effect in this complex scenario requires careful consideration of variables and the implementation of rigorous research in data analysis.
-To estimate the effect of a policy, a common approach is to use partially linear regression as in Robinson (1988), which involves specifying separate regression models for both the treatment and outcome variables.
+Recent methodological advances seek to formalize these adjustments.
+Gaebler et al.
+(2022) argue that many policing studies lack explicit causal models, leading to implicit identifying assumptions.
+Jung et al.
+(2018) propose a risk-adjusted regression framework to mitigate included- and omitted-variable bias and to focus disparity estimates on differences not explained by legitimate police-relevant risk factors.
+Recently, Huang et al.
+(2024) introduce a mobility-adjusted causal estimand, demonstrating that individuals’ movement patterns fundamentally shape exposure to policing.
+Another potential threat to identification in this context is selection bias, since stop and search events often arise because of a form of racial bias.
+The data we get to observe arise from interactions the police have with the civilians they choose to stop, and are certainly not a random sample of all police-civilian interactions.
+To estimate causal effects in such complex settings, applied researchers often adopt flexible semiparametric models.
+A commonly used tool is the partially linear regression framework (Robinson, 1988), which specifies separate regression functions for the treatment and outcome.
 The model for the treatment variable is often termed as ‘propensity score’ model.
 Propensity score adjustments (Rosenbaum and Rubin, 1983) have been extensively used to reduce confounding bias in estimating causal effects.
-The propensity score is defined as the conditional probability of the treatment assignment given confounding covariates.
-The propensity score can be used to break the dependence between confounders and exposure, to create balance in the distribution of confounders across exposure groups, and to facilitate correct inference.
-However, a key statistical challenge in applying partially linear regression for Stop & Search practices is the presence of high-dimensional confounders, where the parameter space for nuisance parameters grows with the sample size.
-In this case, the traditional semi-parameter theory might not offer valid inference for the parameter of interest.
-To overcome this challenge, Chernozhukov et al.
-(2018) proposed the use of machine learning methods to estimate the nuisance parameter and provided a simple and root-$n$ consistent procedure to estimate the parameter of interest via the Neyman moment equation and sample splitting.
-Another challenge arises, when analyzing Stop & Search data, due to the strong prior beliefs and assumptions often held by policymakers and law enforcement agencies.
-These priors may reflect institutional perspectives on crime prevention strategies or operational practices, and we need to account for these prior beliefs in order to ensure robust and objective analysis, fostering evidence-based policy decisions.
-Therefore, a prior-to-posterior Bayesian analysis is preferable.
-The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the effect from a practical perspective.
+The propensity score is defined as the conditional probability of receiving treatment given confounding covariates.
+The propensity score can be used to break the dependence between confounders and exposure, to create balance in the distribution of confounders across exposure groups, and to allow valid causal inference.
+However, a key statistical challenge in applying partially linear regression for stop and search practices is the presence of high-dimensional confounders, where the parameter space for nuisance parameters grows with the sample size.
+In this case, traditional semi-parameter theory might not offer valid inference for the parameter of interest.
+To overcome these challenges, Chernozhukov et al.
+(2018) proposed the use of machine learning methods to estimate the nuisance parameters and provided a simple and root-$n$ consistent procedure to estimate the parameter of interest via the Neyman moment equation and sample splitting.
+Another challenge arises, when analyzing stop and search data, due to the strong prior beliefs and assumptions often held by policymakers and law enforcement agencies.
+These priors may reflect institutional perspectives on crime prevention strategies or operational practices, but might also be potentially biased or based on anecdotal evidence.
+A prior-to-posterior Bayesian analysis allows to incorporate beliefs that are based on prior data-driven evidence, fostering evidence-based policy decisions.
+Further, unlike approaches that rely on large-sample approximations, which may be unreliable in finite samples or when complex, high-dimensional nuisance models are required, Bayesian inference can deliver more stable and valid uncertainty quantification even in small sample sizes that often arise in areal-level analyses.
+In this work, we aim to incorporate Bayesian inference within a double machine learning framework to obtain coherent probabilistic statements about the effects of disproportionality, enhance the stability of estimation, and allow the incorporation of domain knowledge.
+This offers a coherent updating framework for regularization and prior information, which are valuable in high-dimensional or weak-signal settings common in health and social policy applications.
 There is growing interest in the application of Bayesian methodology to two-stage propensity score regression analysis; however, most of existing Bayesian approaches require a full parametric specification for both structural equations in both regression models (see, for example McCandless et al., 2010; Kaplan and Chen, 2012).
@@
 Recently, Luo et al.
-(2023) proposed to draw inference for the posterior from a Bayesian predictive distribution via a Dirichlet process model, extending the Bayesian bootstrap, and opening up the possibility to explore the new direction on doubly robust causal inference based on a non-parametric specification.
+(2023) proposed to draw inference for the posterior from a Bayesian predictive distribution via a Dirichlet process model, extending the Bayesian bootstrap, and opening up the possibility of performing doubly robust causal inference based on a non-parametric specification.
 In their work, the posterior samples are generated by resampling weights from the Pólya urn.
@@
 (2022) proposed to use the Gaussian process (GP) regression, and then the MCMC estimate is plugged into estimating equation.
-However, in present of high-dimensional confouders, those approaches can be slow when implementing Markov chain Monte Carlo (MCMC).
-In this paper, we aim to marry between the Bayesian method and the semi-parametric double machine learning method and provide a computationally efficient procedure to draw inference on the parameter of interest.
+However, despite these advances, computational efficiency and theoretical guarantees remain open challenges, where covariates may be high dimensional and the nuisance components complex.
+In this paper, we aim to marry the Bayesian method and the semi-parametric double machine learning framework and provide a computationally efficient procedure to draw inference on the parameter of interest.
 We argue that our method bridges this gap by demonstrating how the approximate frequentist distribution theory can find an equally effective interpretation within a Bayesian framework in the high-dimensional setting.
 In addition, we show that the proposed procedure generates a valid posterior according to Monahan and Boos (1992), indicating a valid putative ‘posterior’ density computed by a non-standard method should still make probability statements consistent with Bayes’ rule.
-1.3 Plan of paper
 The remainder of this paper is organized as follows.
-In Section 2, we review the estimation procedure of partially linear regression, and introduce the notion of approximating Bayesian formulations and how to perform inference via the Neyman moment equation.
+Section 2 describes the motivating dataset used in this paper.
+In Section 3, we articulate the causal question for the study and define the associated causal estimand.
+In addition, we review the estimation procedure of partially linear regression, and introduce the notion of approximating Bayesian formulations and how to perform inference via the Neyman moment equation.
 We verify the validity of the proposed posterior inference in the spirit of Monahan and Boos (1992) in Section 3.5.
-Section 3 shows some simulation examples which compare the proposed method with some other frequentist and Bayesian approaches, following with Stop & Search data analysis in Section 4.
+Section 4 shows some simulation examples which compare the proposed method with some other frequentist and Bayesian approaches, following with Stop and Search data analysis in Section 5.
 Finally, Section 6 presents some concluding remarks and future research directions.
-2 Statistical methodology
-To estimate the effect of racial dispropotionality, we first consider the partially linear regression in Robinson (1988):
-$Y=\mu(X)+\beta D+U,\;\;\;\mathbb{E}(U|X,D)=0,\;\;\;\mathbb{E}(D|X)=\pi(X)\$ (1)
-where $Y$ is the outcome variable, $D$ is the treatment variable of interest (i.e., the proportion of Black people), and $X=\left(X_{1},\ldots,X_{p}\right)$ is a list of counfounders.
-Given independent and identically distributed data, $\left\{Z_{i}=(Y_{i},D_{i},X_{i}),i=1,\ldots,n\right\}$, the parameter of interest is $\beta$.
-If $D$ is exogenous conditional on $X$, then $\beta$ has the interpretation of the treatment effect parameter.
-Assuming there are no unmeasured confounders Under frequentist inference, the estimand of $\beta$ is consistent if $\mathbb{E}\left[Y\left|X,D\right.\right]$ is correctly specified.
-In addition, one can be constructed via specifying both outcome mean and propensity score models and use them in a combined estimation to yield a debiased estimator.
-To implement such inference, we need to specify the following two models:
-  • Propensity score (PS) model: $\mathbb{E}(D|X)$, and it represents the conditional distribution of treatment given the confounders;
-  • Outcome regression (OR) model: $\mathbb{E}\left(Y\left|X,D\right.\right)$, which represents the conditional distribution of the outcome given the treatment variable and confounder under the observational study.
+2 Motivating example: Racial disproportionality assessment in Stop and Search
+The Stop and Search dataset consists of individual stop and search monthly records in London from January 2019 to December 2023, including date and time, street-level location, ethnicity, gender and age of the person stopped, legislation, object searched and outcome.
+Further description of this dataset can be found in the Appendix.
+Based on 2.6 million stop and search incidents from January 2019 to December 2023 in London aggregated within the borough level, the objective of this study is to examine how the level of disproportionality in stop and search practices for expressive crime targeting Black individuals varies across London boroughs.
+That is, the focus is on borough-level inequality in policing outcomes rather than on the overall presence or absence of racial bias across the city.
+Even if disproportionality is high across London, the analysis seeks to identify whether certain boroughs exhibit relatively greater or lesser disproportionality than the average level in London.
+Therefore, we define the outcome variable as the disproportionality index (DI) for expressive crime targeting Black people in each Borough (total 33 Boroughs) and it can be calculated as:
+[…]
+This index contrasts a Borough-specific stop and search rate involving Black individuals and the average rate in London, where $\text{DI}=1$ indicates no disproportionality, whereas $\text{DI}<1$ indicates a stop and search rate below expectation while $\text{DI}>1$ indicates a stop and search rate above expectation.
+Figure 1: DI for Black People in Stop & Search for Expressive Reasons
+Figure 1 illustrates the DI across boroughs in London.
+The South-West London boroughs of Richmond upon Thames and Kingston upon Thames and the City of London have the largest DI values.
+These boroughs have considerably larger index values, revealing stop and search rates above expectation among the Black population.
+In contrast, east end London boroughs, especially the riverside boroughs — Lewisham, Greenwich, Bexley, Newham, and Barking and Dagenham — all have low DI scores.
+Boroughs nearer the perimeter of the city, like Enfield and Croydon, also have lower DI scores.
+This trend indicates a gradient in the distribution, with more central and wealthy areas tending to have larger index values and outer or more residential boroughs having comparatively lower scores.
+To evaluate the effect of disproportionality, we use the percentage of Black people in each borough as the treatment variable.
+Specifically, areas with a higher percentage of Black residents might experience more police attention, which could inflate the DI.
+If policing practices were biased or influenced by racial profiling, the Stop and Search rates for Black individuals might rise disproportionately, driving up the DI.
+Moreover, in boroughs with larger Black populations, more recorded crimes against Black individuals could naturally occur, but if the rate of police actions exceeds the expected level in London based on population share, it would signal disproportionality.
+For example, if Black residents make up 30% of a borough but account for 60% of Stop and Search incidents, the DI would be elevated, demonstrating a disproportionate impact.
+The post-treatment variable considered in Zhao et al.
+(2022), police detainment, can be viewed as analogous to the outcome in this paper, since the disproportionality index is constructed from the frequency of “detainments” or stop interactions involving individuals from minority backgrounds.
+The key distinction is that our analysis relies on areal-level data, whereas Knox et al.
+(2020); Zhao et al.
+(2022) work with individual-level observations.
+Moreover, in our setting, outcomes for minority groups are not directly contrasted with those for non-minority groups.
+Instead, the potential outcomes for minority individuals correspond to different “doses” of exposure across districts.
+From a policy perspective, this raises the question of whether such patterns can be interpreted as evidence of potential discrimination.
+3 Methodology
+3.1 Notation and causal assumptions
+Let $\left\{Z_{i}=(Y_{i},D_{i},X_{i}),i=1,\ldots,n\right\}$ denote independent and identically distributed data measured on $n$ study units (e.g., different boroughs), where $Y_{i}$ denotes an observed outcome (e.g., the disproportionality index DI), $D_{i}$ represents an exposure or treatment, potentially continuous (e.g., percentage of Black people in an area), and $X_{i}$ is a vector of potential confounders measured on unit $i$.
+The primary focus of our inference is on average potential outcomes (APOs) for a population under a hypothetical exposure.
+We let $Y_{i}(d)$ be the (potential) DI that would be observed in borough $i$ if the proportion of the Black population were set to $d$.
+The target estimand is the average causal effect of the Black population proportion on stop and search disproportionality, $\beta=\mathbb{E}\left[\frac{\partial Y(d)}{\partial d}\right]$, capturing the average difference in disproportionality.
+This estimand captures the causal effect of borough-level racial composition on the degree of disproportionality in stop and search incidents involving Black individuals for expressive crimes.
+Given these confounders, denoted as $X_{i}$, in borough $i$, our framework relies on the usual core causal assumptions: consistency, Stable Unit Treatment Value (SUTVA), no unmeasured confounding (NUC), and positivity (Antonelli et al., 2022; Li et al., 2023).
+Under these assumptions, the distribution of potential outcomes is identified from observed data via the g-formula: $E\left[Y(d)\right]=\mathbb{E}\left\{\mathbb{E}\left[Y|D=d,X\right]\right\}$.
+In the next section, we will specifically focus on partially linear regression to estimate this causal effect using a Bayesian double machine learning approach.
+3.2 Partially linear regression
+To estimate the effect of policy, i.e. racial dispropotionality, we consider the partially linear regression described in Robinson (1988):
+$Y=\mu(X)+\beta D+U,\;\;\;\mathbb{E}(U|X,D)=0,\;\;\;\mathbb{E}(D|X)=\pi(X)\\$ (1)
+where $Y$ is the outcome variable, $D$ is the treatment variable, $X=\left(X_{1},\ldots,X_{p}\right)$ is a list of counfounders and $U$ represents the unobserved error term.
+[…]
+  • Propensity score (PS) model: $\mathbb{E}(D|X)$, which represents the conditional distribution of treatment given the confounders;
+[…]
 As noted in Lee (2018), The model in (1) can be rewritten as a new regression model:
-$Y-\mathbb{E}[Y|\pi(X)]=\beta[D-\pi(X)]+V,\;\;\;V=U+\mu(X)-\mathbb{E}[\mu(X)|\pi(X)]=0,\;\;\;V\perp\!\!\!\perp D|\pi(X)\$ (2)
+$Y-\mathbb{E}[Y|\pi(X)]=\beta[D-\pi(X)]+V,\;\;\;V=U+\mu(X)-\mathbb{E}[\mu(X)|\pi(X)]=0,\;\;\;V\perp\!\!\!\perp D|\pi(X)\\$ (2)
 and therefore $\mathbb{E}(V|\pi(X))=0$.
-The treatment effect parameter, $\beta$, can be estimated via the least square estimation procedure.
-This is equivalent to the Neyman-orthogonal moment equation noted in Chernozhukov et al.
-(2018), and $\beta$, is the solution $\mathbb{E}[\psi(Z;\beta)]=0$, where
+This implies that $cov(D-\pi(X),V)=0$, and the treatment effect parameter, $\beta$, can be estimated via the least square estimation procedure from regressing $Y-\mathbb{E}[Y|\pi(X)]$ on $D-\pi(X)$.
+However, Hahn (1998) demonstrated that the estimator is not semiparametrically efficient when both $\mathbb{E}[Y|\pi(X)]$ and $\pi(X)$ are estimated non-parametrically.
+To address this, Chernozhukov et al.
+(2018) proposed estimating $\beta$ using the Neyman-orthogonal moment equation, which achieves the semi-parametric efficiency bound as it protects the estimator against first-order bias from nuisance function estimation errors.
+That is, $\beta$ is the solution $\mathbb{E}[\psi(Z;\beta)]=0$, where
 $\psi(Z;\beta)=[D-\pi(X)][Y-\beta D-\mu(X)].$
+We can estimate nuisance functions, $\pi(X)$ and $\mu(X)$, non-parametrically or using machine learning (ML) methods regardless of the dimension of $X$.
 In this way, we can reduce the high-dimensional problem to a single parameter problem if we assume $D$ is a scalar.
 This approach requires that we have access to both $\mu(X)$ and $\pi(X)$.
-The challenge arises in the high-dimensional setting, i.e., $p>n$; however, both of these functions can be estimated via machine learning (ML) methods.
+The challenge arises in the high-dimensional setting, i.e., $p>n$; however, both of these functions can be estimated via ML methods.
 This double ML specification gives us flexibility to specify the structures of $\mu(X)$ and $\pi(X)$ while focusing on the treatment effect.
-However, it would be difficult to perform conventional Bayesian analysis with an strong prior input from the policy makers as there is in no distributional form for the conditional outcome mean and PS models.
+However, it would be difficult to perform conventional Bayesian analysis with an strong prior input from the policy makers as there is no distribution assumption in this semi-parametric procedure.
 Conventional Bayesian inference focuses on updating prior belief in light of the data, and the data are summarized in the form of the likelihood.
@@
 In the next section, we will seek solutions to incorporate the Neyman-orthogonal approach into a fully Bayesian procedure and demonstrate how the approximate frequentist distribution theory, which rests on moment constraint assumptions about distributions, can find an equally effective interpretation within a Bayesian framework.
-2.1 Bayesian non-parametric procedure for the Neyman moment equation
-[…]
-[…]
+3.3 Bayesian non-parametric procedure for the double machine learning method
+Suppose we assume that there is a set of Neyman-orthogonal equations $\psi(Z;\beta)$ such that $\mathbb{E}[\psi(Z;\beta)]=0$, for all $\beta$.
+The objective is to find a non-parametric approximate likelihood, $p(z)$, to the true data generating model $f(z|\beta)$.
+Therefore, we can define some discrepancy $\delta(f|p)$, subject to $\int\psi(z;\beta)p(z)dz=0$ and $\int p(z)dz=1$.
+This becomes an optimization problem:
+$\min_{p}\delta(f|p)\text{ subject to }\int\psi(z;\beta)p(z)dz=0\;\;\int p(z)dz=1,\;\;\forall\beta\in\mathbb{R}.$ (3)
 If we specify $p$ as nonparametric, and use $\delta(f|p_{1},\ldots,p_{n})$ to measure the discrepancy between the true data generating model and the approximating model.
@@
 where $\lambda$ is a user-specified parameter. From Baggerly (1998), this function can be rewritten as
-$\text{CR}(p)=\left\{\begin{array}[]{cc}-2\sum_{i=1}^{n}\log(np_{i}),&\lambda=0\\ &\\ 2n\sum_{i=1}^{n}p_{i}\log(np_{i}),&\lambda=-1\\ &\\ \frac{2}{\lambda(1+\lambda)}\sum_{i=1}^{n}\left[(np_{i})^{-\lambda}-1\right],&\lambda\neq-1\text{ or }0.\\ \end{array}\right.$
-Therefore, the solution in (4) is the nonparametric likelihood which seeks to reweight the sample so that it can also satisfy the moment condition (Qin and Lawless, 1994).
-It has been proven to possess many properties of the conventional parametric likelihood theory (Owen, 2001).
-A class of generalized empirical likelihood functions are studied in Imbens et al.
-(1998); Chernozhukov and Hong (2003); Newey and Smith (2004).
-In our example, we want to place a prior on the treatment effect parameter, $\beta$, directly, and update this prior in light of the data observed.
-Therefore, we can replace $f\left(z_{i}\left|\beta\right.\right)$ with the profile likelihood, $p_{i}$, and then the posterior distribution for $\beta$ becomes
+[…]
+[…]
 $\pi\left(\beta\left|\mathbf{z}\right.\right)\propto\pi_{0}\left(\beta\right)\prod\limits_{i=1}^{n}p_{i}.$
@@
 where $\kappa_{1}$ and $\kappa_{2}$ are the Lagrange multipliers. Setting $\partial\text{GEL}/\partial p_{i}=0$ yields the extreme of the form
-$p_{i}=\left\{\begin{array}[]{cc}\frac{1}{n}\left[1+s+t\psi(z_{i};\beta)\right]^{-1/(1+\lambda)},&\lambda\neq-1\\ &\\ s\exp\left[t\psi(z_{i};\beta)\right],&\lambda=-1\\ \end{array}\right.$ (5)
+[…]
 where $s$ and $t$ are normalized constant and determined by
@@
 Therefore, by substituting (5) into the posterior distribution, we obtain the following posterior distribution
-$\pi(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\left\{\begin{array}[]{cc}\prod_{i=1}^{n}\frac{1}{n}\left[1+\hat{s}(z,\beta)+\hat{t}(z,\beta)\psi(z_{i};\beta)\right]^{-1/(1+\lambda)}&\lambda\neq-1\\ &\\ \prod_{i=1}^{n}\frac{\exp\left(\hat{t}(z,\beta)^{\top}\psi(z_{i};\beta)\right)}{\sum_{j=1}^{n}\exp\left(\hat{t}(z,\beta)\psi(z_{j};\beta)\right)}&\lambda=-1\\ \end{array}.\right.$ (6)
+$\pi(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\left\{\begin{array}[]{cc}\prod_{i=1}^{n}\frac{1}{n}\left[1+\hat{s}(z,\beta)+\hat{t}(z,\beta)\psi(z_{i};\beta)\right]^{-1/(1+\lambda)}&\lambda\neq-1\\ &\\ \prod_{i=1}^{n}\frac{\exp\left(\hat{t}(z,\beta)^{\top}\psi(z_{i};\beta)\right)}{\sum_{j=1}^{n}\exp\left(\hat{t}(z,\beta)\psi(z_{j};\beta)\right)}&\lambda=-1\\ \end{array}.\right$.
+(6)
 The following algorithm summarizes the computation step to the Bayesian generalized empirical likelihood method.
 Algorithm 1 Algorithm to obtain the posterior sample of $\beta$ via the Bayesian generalized empirical likelihood.
-1:$\mathcal{D}=(z_{1},\ldots,z_{n})$
+[…]
 2:Estimate $\pi(x)$ and $\mu(x)$ using some ML methods.
 3:for $j$ to $1:J$ do
-4:     Sample $\beta^{(j)}\sim\pi_{j}(\beta\left|\mathbf{z}\right.)$ using the MCMC approach, where $\pi_{j}(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\prod_{i=1}^{n}p_{i}^{(j)}$.
+[…]
 5:  $\left(p^{(j)}_{1},\ldots,p^{(j)}_{n}\right)$ is the solution to $\min_{p_{1},\ldots,p_{n}}\text{CR}(p)\text{ subject to }\sum_{i=1}^{n}p_{i}=1,\;\;p_{i}\geq 0,\;\;\sum_{i=1}^{n}p_{i}\psi(z_{i};\beta^{(j-1)})=0$
@@
 3.4 The role of sample splitting procedures
-In Chernozhukov et al.
-(2018), they found that sample splitting plays a key role in reducing the bias when estimating $\beta$.
-The remainder term in the Taylor expansion involves the product between the error term in the PS model and the bias using ML method to estimate $\mu(\cdot)$.
+When using ML methods to estimate the nuisance functions, Chernozhukov et al.
+(2018) found that sample splitting plays a key role in reducing the bias when estimating $\beta$.
+When showing the consistency, the remainder term in the Taylor expansion involves the product between the error term in the PS model and the bias using ML method to estimate $\mu(\cdot)$.
 In some cases, this remainder term might not vanish when $n\to\infty$ as the two terms are correlated.
@@
 In this simple example, both methods indicate some biases from the full sample approach while the Bayesian method has a slightly smaller bias.
-In light of this consideration, Chernozhukov et al.
+To overcome this issue, Chernozhukov et al.
 (2018) proposed the use of sample splitting, that is, the data are partitioned into $K$ groups.
 The functions $\hat{\mu}_{k}(\cdot)$ and $\hat{\pi}_{k}(\cdot)$ are estimated using all the data excluding the $k$th group.
-Then the double ML estimator for $\beta$ is the solution to $01/K\sum_{k=1}^{K}\mathbb{E}_{k}[\psi(Z;\beta)]=$ 0, where $\mathbb{E}_{k}(\cdot)$ is the empirical expectation over the $k$th fold of the data.
+Then the double ML estimator for $\beta$ is the solution to $1/K\sum_{k=1}^{K}\mathbb{E}_{k}[\psi(Z;\beta)]=0$, where $\mathbb{E}_{k}(\cdot)$ is the empirical expectation over the $k$th fold of the data.
 This creates the independence between two terms, leading to unbiased estimation.
@@
 Algorithm 2 Algorithm to obtain the posterior sample of $\beta$ via the Bayesian generalized empirical likelihood with sampling splitting.
-1:$\mathcal{D}=(z_{1},\ldots,z_{n})$
+[…]
 2:Partition the data into $K$ groups (roughly equal size), $\mathcal{D}=\left\{\mathcal{D}_{1},\ldots,\mathcal{D}_{K}\right\}$.
@@
 7:  end for
-8:     Sample $\beta^{(j)}\sim\pi_{j}(\beta\left|\mathbf{z}\right.)$ using the MCMC approach, where $\pi_{j}(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\prod_{i=1}^{n}p_{i}^{(j)}$.
+[…]
 9:end for
@@
 The second column of Figure 2 displays the results using two-fold sample splitting, i.e., $K=2$.
-Both the double machine learning (ML) methods and the amended Bayesian approach effectively eliminate bias from the full-sample estimation procedure.
+Both the double ML methods and the amended Bayesian approach effectively eliminate bias from the full-sample estimation procedure.
 The posterior distribution obtained using Algorithm 2 exhibits similar distributional properties across 10,000 replications.
@@
 As demonstrated in Chernozhukov et al.
-(2018), if the estimated functions of $\mu(x)$ and $\pi(x)$ converge to the true functions in probability under some regularity conditions, the estimator for $\beta$ in (2) will converge in distribution to a normal distribution, centered at $\beta$ with variance, $\mathbb{E}[(D-\pi(X))^{2}]^{-1}\mathbb{E}[(Y-\beta D-\mu(X))^{2}]$.
+(2018), if the estimated functions of $\mu(\cdot)$ and $\pi(\cdot)$ converge to the true functions in probability under some regularity conditions, the estimator for $\beta$ in (2) will converge in distribution to a normal distribution, centered at $\beta$ with variance, $\mathbb{E}[(D-\pi(X))^{2}]^{-1}\mathbb{E}[(Y-\beta D-\mu(X))^{2}]$.
 This is the semi-parametric efficiency bound for $\beta$.
-The ‘posterior’ density in Algorithm 2, $\pi(\beta|z_{1:n})$, is a non-standard posterior inference as the likelihood is replaced by the profile likelihood with plug-in estimates for $\mu(x)$ and $\pi(x)$.
-Despite the theoretical support of the ETEL case in Schennach (2007); Chib et al.
-(2018), one would want to assess if this plug-in coupled with sample-splitting strategy is valid Bayesian inference.
-That is, if interval $S_{\alpha}(z)$ is a designated $1-\alpha$ probability interval, a ‘posterior’ density in Algorithm 2, $\pi(\beta|z_{1:n})$, will have the property $\mathbb{P}_{\pi}\left[\beta\in S_{\alpha}(z))\right]=1-\alpha$, if $Z_{1},\ldots,Z_{n}$ are drawn from the true data generating model.
-Monahan and Boos (1992) proposed a notion of proper Bayesian inference by replacing the parametric likelihood with an alternative likelihood function.
-They stated that the ‘posterior’ density, which derives from the alternative likelihood, should follow the law of the probability deriving from the Bayes’s rule.
-The posterior density is defined as valid by coverage if $\mathbb{P}_{\pi}\left[\beta\in S_{\alpha}(z))\right]=1-\alpha$, if $Z_{1},\ldots,Z_{n}$ are drawn from the true data generating model.
-The posterior coverage set, $S_{\alpha}(z)$, resulting from a valid posterior, should achieve nominal coverage under the joint measure of $Z$ and $\theta$, that is, $P_{\pi}(\beta\in S_{\alpha}(z))$ should have expectation $1-\alpha$ for data generated under the measure $\pi_{0}(\beta)f(z|\beta)$ on $(\beta,Z)$ for every absolutely continuous prior, $\pi_{0}(\ldotp)$.
-To verify this property, let
 […]
-and if $\pi(\beta|\mathbf{z})$ is a valid posterior, then $H$ follows Uniform$(0,1)$.
-In practice, if we generate $\beta_{k}(k=1,\ldots,m)\sim\pi_{0}(\ldotp)$ and the data, $z_{1:n}^{(k)}$, from $f(\ldotp|\beta_{k})$, and compute the posterior according to Algorithm 2.
-Then we can obtain $H_{k}$ based on (7) by replacing $\beta$ with $\beta_{k}$.
-If the distribution of $H_{k}$ follows the uniform distribution, then the posterior distribution generated from Algorithm 2 is defined as a coverage proper posterior, yielding valid posterior inference.
-This evaluation method for the validity of posterior inference has also been applied to verify the correctness of Bayesian computation (see for example, Talts et al., 2018).
+$H=\int_{-\infty}^{\beta}\pi(\varphi|z)d\varphi,$ (7)
+[…]
 In light of this approach, we investigate the validity of the proposed generalized empirical likelihood approach via a simulation study using the same set-up with the sample-splitting simulation, with each $\beta_{k}$ ($k=1,\ldots,10000)$ generating from $N(1,2)$.
@@
 For all cases, it suggests that simulation $H$ values follow uniform distribution and indicates proper posterior inference according to Monahan and Boos (1992) when $\lambda=0,-1,-1/2$.
-Therefore, this simulation result gives us evidence that the proposed method according to Algorithm 2 can be regarded as a valid approach for posterior inference in terms of coverage.
-Figure 2: Histgrams for $H$ statistics using the Bayesian generalized empirical likelihood. $P$-values represents the Kolmogorov–Smirnov test for uniformity of $H$.
+Therefore, this simulation result gives us evidence that the proposed method according to Algorithm 2 can be regarded as a valid approach for posterior inference.
+Figure 3: Histograms for $H$ statistics using the Bayesian generalized empirical likelihood. $P$-values represents the Kolmogorov–Smirnov test for uniformity of $H$.
 4 Simulation
@@
   • BDR-HD: The Bayesian doubly robust high-dimension method proposed in Antonelli et al.
-(2022), where the propensity score and outcome are estimated via regression models with the Gaussian process prior, and then the MCMC estimate is plugged in to a doubly robust estimator.
+(2022), where the propensity score and outcome are estimated via regression models with the GP prior, and then the MCMC estimate is plugged in to a doubly robust estimator.
 The variance is adjusted through the frequentist bootstrap so that it will achieve the nominal coverage rate.
@@
 For all the Bayesian methods, we generate $5,000$ MCMC samples and $1,000$ burn-in iterations each simulation with $1,000$ simulation replicates.
+The code for the simulation is publicly available on our GitHub page at https://github.com/yumcgill/Bayesian-DML.
 4.1 Binary exposure
-In this case, we consider $D$ is binary and simulate $X=(X_{1},X_{2},\ldots,X_{p})\sim\mathcal{N}_{p}\left(0,\Sigma\right)$ and $\Sigma_{ij}=1$ if $i=j$ and 0.3 otherwise, and then simulate
-[…]
+In this case, we consider $D$ is binary and simulate
+$X=(X_{1},X_{2},\ldots,X_{p})\sim\mathcal{N}_{p}\left(0,\Sigma\right),\Sigma_{ij}=\begin{cases}1,&\text{if }i=j,\\ 0.3,&\text{if }i\neq j,\end{cases}$
+and then simulate
+$D\left|X\right$.  $\sim\text{ Bernoulli}\left(\text{expit}\left(0.3X_{1}+0.2X_{2}-0.4X_{5}\right)\right)$
 $Y\left|D,X\right$. $\sim\mathcal{N}\left(D+0.5X_{1}+X_{3}-0.1X_{4}-0.2X_{7},1\right).$
@@
 Table 1 shows the performance of ETEL, EL, HD, DML across machine learning approaches (Lasso, Random Forest, and Neural Network) and BDR-HD in terms of bias, root mean squared error (RMSE), and coverage rate.
-For the proposed Bayesian method, it maintain low bias and RMSE with relatively high coverage rates.
+For the proposed Bayesian method, it maintains low bias and RMSE with relatively high coverage rates.
 Specifically, EL and HD generally achieve slightly better precision than ETEL, with HD sometimes offering the lowest bias.
 In terms of machine learning approaches, neural networks, although effective in reducing bias, show higher RMSE and lower coverage rates compared to Lasso and Random Forest.
-DML method demonstrates quite similar results with the proposed Bayesian method, in line with our previous demonstration about the uncertainty quantification.
+The DML method demonstrates quite similar results with the proposed Bayesian method, in line with our previous demonstration about the uncertainty quantification.
 BDR-HD shows similar bias as other methods, but it achieves the lowest RMSE and the highest coverage rate due to the extra bootstrap step to adjust the posterior variance.
@@
 In this example, the coefficient associated with $D$ is the main regression coefficient and is the parameter of interest.
-We set $p=40$ and $n=40$, and in the proposed Bayesian method, we assign an informative prior, $\mathcal{N}(1,2)$, to $\beta$.
+We set $p=40$ and $n=40$, and in the proposed Bayesian method, we assign a relatively informative prior, $\mathcal{N}(1,2)$, to $\beta$.
 Table 2 shows the results of bias, RMSE and coverage rate across all methods.
-Most of the methods perform well in terms of bias, ranging between $-0.01$ and $0.06$, while the RMSE varies slightly, with most values in between $0.13$ and $0.19$.
-In particular, Lasso-based approaches have very low bias ($0.01-0.02$), while random forest based methods show slightly higher bias.
-In terms of RMSE, the proposed Bayesian method displays smallest RMSEs $(0.13-0.14)$ across all ML methods.
-DML methods showing a slightly higher RMSE $(0.16-0.19)$, while BDR-HD has a much higher RMSE.
-Regarding coverage rates, EL-based methods always achieve higher rates $(90.6-94.0)$, while ETEL-based methods show a bit lower coverage $(85.8-89.9)$.
+Most of the methods perform well in terms of bias, ranging between 0.01-0.01 and $0.06$, while the RMSE varies slightly, with most values in between $0.13$ and $0.19$.
+In particular, Lasso-based approaches have very low bias (0.010.020.01-0.02), while random forest based methods show slightly higher bias.
+In terms of RMSE, the proposed Bayesian method displays smallest RMSEs (0.130.14)(0.13-0.14) across all ML methods.
+DML methods showing a slightly higher RMSE (0.160.19)(0.16-0.19), while BDR-HD has a much higher RMSE.
+Regarding coverage rates, EL-based methods always achieve close to nominal level rates (90.694.0)(90.6-94.0), while ETEL-based methods show a bit lower coverage (85.889.9)(85.8-89.9).
 BDR-HD reaches the coverage rate at 95.9, closest to the nominal level.
 We also notice that methods using neural network generally have lower coverage rates than methods using Lasso and random forest.
-Results indicate that the proposed Bayesian method, coupled with an informative prior, performs well overall in terms of balancing low bias, low RMSE, and adequate coverage rates.
+Results indicate that the proposed Bayesian method with EL-based likelihood, coupled with an informative prior, performs well overall in terms of balancing low bias, low RMSE, and adequate coverage rates.
 Table 2: Continuous exposure: Simulation results of the marginal causal effect under high-dimensional settings, with true value equal to 1, on 1000 simulation runs with $n=40$ and $p=40$.
@@
 BDR-HD 0.03 0.75 95.9
-4 Application: Stop and Search in London, UK
-4.1 Stop and Search data
-In this section, we apply the proposed methodology to the Stop & Search data in London to enhance our understanding of the impact of the racial disproportionality, particularly on expressive crimes.
-The Stop & Search dataset consists of individual stop and search monthly records in London from January 2019 to December 2023, including date and time, street-level location, ethnicity, gender and age of the person stopped, legislation, object searched and outcome.
-The data have been sourced from the Metropolitan Police and City of London police force and published in the Single Online Home National Digital Team under Open Government Licence v3.0.
-They are available openly from https://data.police.uk/docs/method/stops-at-location/.
-Before publishing, the location coordinates of the stop are anonymized (detailed methods of anonymization from the publisher can be found at https://data.police.uk/about/#location-anonymisation and the age of the person stopped has been adjusted to a corresponding age group (e.g.
-18-24).
-Category Searched Objects Legislation
-Acquisitive objects ‘Stolen goods’, ‘Article for use in theft’ ‘Police and Criminal Evidence Act 1984 (section 1)’, ‘Criminal Justice Act 1988 (section 139B)’, ‘Police and Criminal Evidence Act 1984 (section 6)’
-Expressive objects ‘Offensive weapons’, ‘Anything to threaten or harm anyone’, ‘Firearms’, ‘Fireworks’, ‘Evidence of offences under the Act’, ‘Crossbows’, ‘Game or poaching equipment’ ‘Police and Criminal Evidence Act 1984 (section 1)’, ’Criminal Justice Act 1988 (section 139B)’, ‘Police and Criminal Evidence Act 1984 (section 6)’, ‘Criminal Justice and Public Order Act 1994 (section 60)’, ‘Firearms Act 1968 (section 47)’
-Drug objects ‘Controlled drugs’, ‘Psychoactive substances’ ‘Misuse of Drugs Act 1971 (section 23)’, ‘Psychoactive Substances Act 2016 (s36(2))’
-In this research, 2.6 million Stop & Search incidents have been compiled for the observation period from January 2019 to December 2023, on seven legislative grounds (Legislation in the above table) for 11 types of Searched Objects.
-Upon the aggregation of the data into three listed categories, drug objects related stop and search incidents took up 64% over the observation period, as compared to 12% for acquisitive objects and 24% for expressive objects.
-The searched subjects’ ethnic information have been collected from “officer defined ethnic group”.
-4.2 Demographic data on ethnic groups
-The demographic data are taken from Census 2021 data by the Office for National Statistics (ONS) and are available from https://www.nomisweb.co.uk/sources/census_2021.
-The demographic data features have been collocated at local authority, “borough”, level as the analytical geographical unit in this research.
-The ethnic groups in Census 2021 are mainly 5 groups including “Asian, Asian British, Asian Welsh”, “Black, Black British, Black Welsh, Caribbean or African”, “Mixed or Multiple”, “White”, and ”Other ethnic group” (ONS, 2021).
-To be consistent with the categories of ethnic groups in the Stop & Search dataset, we further group the ethnic groups into four groups as: “Asian, Asian British, Asian Welsh”, “Black, Black British, Black Welsh, Caribbean or African”, “White”, and “Others” (including Mixed and Other groups in the official Census dataset).
-From the latest Census 2021 data, ethnic groups’ proportions in London are unevenly distributed in that, White people take up 53.73%, followed by Asian people at 20.7%, Black people at 13.52%, and Others at 12.05%.
-Exploratory data analysis on Stop and Search subjects’ demographics, especially the ethnic compositions, has found disproportionally distributed Stop & Search among the ethnic groups.
-White people took up on average 40% over time (ranged from 38% to 42%), Black people took up on average 38% during the observation period (ranged from 36% to 40%), Asian people took up on average 17% (ranged from 14% to 18%), then Others at 5% (ranged from 4% to 6%), as depicted in Figure 1.
-Figure 3: Stop and Search Subjects Ethnic Groups Proportions
-4.3 Outcome definition
-To apply our method, we define the outcome variable as the dispropotionality index (DI) for expression crime targeting Black people in each Borough (total 33 Boroughs) and it can be calculated as:
-$\text{DI}_{i}=\frac{\text{rate of black people in Stop \& Search for expression crime in Borough }i}{\text{rate of average black people in Stop \& Search for expression crime in London}}.$
-This index compares the rate of Black people in the Stop & Search for expression crime in each Borough to the average rate in London:
-  • If $\text{DI}=1$, it indicates no disproportionality;
-  • If $\text{DI}<1$, it indicates too little Stop & Search below expectation;
-  • If $\text{DI}>1$, it indicates too much Stop & Search above expectation.
-Figure 4: DI for Black People in Stop and Search for Expressive Reasons
-Boroughs in South-West London, such as Richmond upon Thames and Kingston upon Thames, as well as City of London, are depicted in dark red with highest DI; whilst boroughs in the east alongside the banks, such as Lewisham, Greenwich, Bexley, Newham, Barking and Dagenham, as well as the very North Enfield and far South Croydon, have presented lower DI among boroughs.
-Figure 5: Monthly Stop and Search in London (2019-2023)
-4.4 Results
-We use the percentage of black people in each borough as the treatment variable.
-The estimated effect is adjusted by variables such as percentage of males (gender), unemployment number and migrant rate, with a total of 31 variables.
+5 Real data application
+In this section, we apply the proposed methodology to the Stop and Search data, described in Section 2 in London to enhance our understanding of the impact of the racial disproportionality, particularly on expressive crimes.
+To conduct an informative analysis, we assign an informative prior to the model, specifically $\mathcal{N}(0,2)$, which reflects our initial assumption that there is no effect of disproportionality in stop and search practices.
+In terms of confounders, we include variables such as percentage of males (gender), unemployment number and migrant rate, with a total of 31 variables.
 A summary of all variables, including their descriptive statistics aggregated in the borough level in the analysis, is given in Table 3.
-We consider the Bayesian partially linear structural equation models with an informative prior, $\mathcal{N}(0,1)$, indicating the prior belief and assumption on no racial disproportion in Stop & Search practice in London.
-We implement the Bayesian generalized empirical likelihood with $\lambda=0,-1,-1/2$ coupled with Lasso, random forest and neural network methods, and generate in total 10000 posterior samples in each case with 1000 burn-in iterations.
+We estimate the both $\pi(\cdot)$ and $\mu(\cdot)$ using all the confounders listed in Table 3, and then apply these estimates in the Neyman-orthogonal equation via the Bayesian generalized empirical likelihood.
+We implement the Bayesian generalized empirical likelihood with $\lambda=0,-1,-1/2$ coupled with Lasso, random forest and neural network methods, and generate in total 10,000 posterior samples in each case with 1,000 burn-in iterations.
 Table 4 shows the results from applying the proposed method.
 Across all cases, the posterior mean estimates are consistently negative, ranging from -0.55 to -0.31.
-In addition, while it shows negligible differences across ETEL, EL and HD, the results for different machine learning methods vary, especially the width of the credible intervals.
-Those for Lasso are the widest, and this might be due to the high correlation among some features, leading to high variability in cross-validation when selecting the optimal tuning parameter.
+In addition, while it shows negligible differences across ETEL, EL and HD, the results for different ML methods vary, especially the width of the credible intervals.
+Those for Lasso are the widest, and this is due to the high correlation among some features, leading to high variability in cross-validation when selecting the optimal tuning parameter.
 Random forest and neural network models have relatively narrower confidence intervals, with the narrowest under EL and HD methods among them.
-Overall, all methods suggest a negative relationship in terms of posterior mean, indicating that disproportionality in Stop & Search aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on Stop & Search against Black people for expressive crimes.
-That is, the higher the percentage of black residents composition in the borough, the lower the predicted value of DI of Stop & Search on Black people for expressive reasons.
-Table 3: Descriptive statistics of the Stop & Search data in London
+Overall, all methods suggest a negative relationship in terms of the posterior mean, indicating that accounting for the proportion of the Black population can help alleviate disproportionality in stop and search practices targeted at the Black community.
+Specifically, when focusing on stop and search incidents involving Black individuals for expressive crimes, we observe that as the percentage of Black residents in a borough increases, the predicted value of the DI for these practices decreases.
+This finding suggests that boroughs with a higher percentage of its population composed of Black individuals will report lower levels of DI in stop and search practices.
+Table 3: Descriptive statistics of the Stop and Search data in London
 Variable Min. Max. Mean Std. Dev.
@@
 Density of Schools 0.85 8.00 2.98 1.83
-Table 4: Posterior mean of the effect of for expression crime targeting black population with associated 95% credible interval.
+Table 4: Posterior mean of the effect of for expressive crime involving black population with associated 95% credible interval.
 Lasso Random forest Neural network
-ETEL $-0.59$ $-0.31$ $-0.41$
-$(-2.17,0.95)$ $(-0.52,-0.14)$ $(-0.73,-0.16)$
-EL $-0.56$ $-0.31$ $-0.40$
-$(-2.21,0.99)$ $(-0.53,-0.14)$ $(-0.78,-0.14)$
-HD $-0.55$ $-0.31$ $-0.41$
-$(-2.18,1.00)$ $(-0.55,-0.14)$ $(-0.78,-0.16)$
+ETEL 0.59-0.59 0.31-0.31 0.41-0.41
+(2.17,0.95)(-2.17,0.95) (0.52,0.14)(-0.52,-0.14) (0.73,0.16)(-0.73,-0.16)
+EL 0.56-0.56 0.31-0.31 0.40-0.40
+(2.21,0.99)(-2.21,0.99) (0.53,0.14)(-0.53,-0.14) (0.78,0.14)(-0.78,-0.14)
+HD 0.55-0.55 0.31-0.31 0.41-0.41
+(2.18,1.00)(-2.18,1.00) (0.55,0.14)(-0.55,-0.14) (0.78,0.16)(-0.78,-0.16)
 6 Discussion
@@
 Our approach also accommodates flexible machine learning models for treatment and outcome, mitigating the impact of model misspecification.
-In our application, we estimate the effect of impact of ethnic disproportionality in Stop & Search on expressive crimes in London with an informative prior, and conclude that: Stop & Search practices in London are disproportionally distributed among London boroughs and disproportionality targeting the Black community is mitigated by considering the proportion of the Black population for expressive crimes.
+In our application, we estimate the effect of impact of ethnic disproportionality in stop and search for expressive crimes in London with an informative prior, and conclude that stop and search practices in London are disproportionally distributed among London boroughs and disproportionality involving the Black community is mitigated by considering the proportion of the Black population for expressive crimes.
 The ethnic composition of demographics plays a significant role in affecting the identified disproportionality, with lower DI alleviated by higher percentage of black population in target boroughs.
+In other words, our analysis reveals that police are more aggressively policing Black individuals in areas that are predominantly white.
+The estimand considered in our application differs from previous work such as that of Zhao et al.
+(2022), in which it does not contrast outcomes from minority groups with outcomes from non-minority groups.
+Also, the outcome of interest in this analysis is the event of a stop interaction, as opposed to the use of police force following a stop interaction or detainment as in Knox et al.
+(2020) and Zhao et al.
+(2022).
+Our DI compares the intensity of stop interactions involving Black individuals across different boroughs of London.
+In this sense, we are not examining whether Black individuals are treated worse than those of other races, but rather we are studying how Black individuals are treated in the different boroughs to detect geographic inequality.
+It could be that racial bias is severe in terms of how minority groups are treated compared to non-minority groups, but that it is done equally across the city.
+A future analysis involving individual-level data comprising different ethnic groups could entail estimating the estimands discussed in Zhao et al.
+(2022) using the proposed Bayesian double machine learning methodology.
+Moreover, we should note that residual spatial autocorrelation was not accounted for in this application.
+An important future extension would involve incorporating a Gaussian random Markov field component to the model to account for spatial dependence.
 Bayesian methods hold significant interest in applied scientific causal research.
@@
 For example, the proposed methodology can be widely applied in other causal settings when the traditional Bayesian set-up requires over-specifying the model condition.
-Moreover, this Bayesian method, coupled with machine learning methods, also presents an important direction for advancement to accommodate complex data structures, such as time-varying treatments, longitudinal data, or hierarchical frameworks in high-dimensional scenarios.
+Moreover, this Bayesian method, coupled with machine learning methods, also presents an important direction for advancement to accommodate complex data structures, such as time-varying treatments, longitudinal data, interference, or hierarchical frameworks in high-dimensional scenarios.
 Acknowledgments
@@
 The authors would like to thank the Evidence and Insight team from MOPAC (Mayor’s Office for Policing and Crime, Greater London Authority) for their valuable comments and suggestions that contributed to the progress of this research.
+Appendix A Additional details regarding the Stop & Search dataset
+A.1 Demographic data on ethnic groups
+The demographic data are taken from Census 2021 data by the Office for National Statistics (ONS) and are available from https://www.nomisweb.co.uk/sources/census_2021.
+The demographic data features have been collocated at local authority, “borough”, level as the analytical geographical unit in this research.
+The ethnic groups in Census 2021 are mainly 5 groups including “Asian, Asian British, Asian Welsh”, “Black, Black British, Black Welsh, Caribbean or African”, “Mixed or Multiple”, “White”, and ”Other ethnic group” (ONS, 2021).
+To be consistent with the categories of ethnic groups in the Stop & Search dataset, we further group the ethnic groups into four groups as: “Asian, Asian British, Asian Welsh”, “Black, Black British, Black Welsh, Caribbean or African”, “White”, and “Others” (including Mixed and Other groups in the official Census dataset).
+From the latest Census 2021 data, ethnic groups’ proportions in London are unevenly distributed in that, White people take up 53.73%, followed by Asian people at 20.7%, Black people at 13.52%, and Others at 12.05%.
+Exploratory data analysis on Stop & Search subjects’ demographics, especially the ethnic compositions, has found disproportionally distributed stop and search events among the ethnic groups.
+White people took up on average 40% over time (ranged from 38% to 42%), Black people took up on average 38% during the observation period (ranged from 36% to 40%), Asian people took up on average 17% (ranged from 14% to 18%), then Others at 5% (ranged from 4% to 6%).
+Figure 4: Stop & Search Subjects Ethnic Groups Proportions
+A.2 Data aggregation procedures
+The data have been sourced from the Metropolitan Police and City of London police force and published in the Single Online Home National Digital Team under Open Government Licence v3.0.
+They are publicly available from https://data.police.uk/docs/method/stops-at-location/.
+Before publishing, the location coordinates of for each stop are anonymized (detailed methods of anonymization from the publisher can be found at https://data.police.uk/about/#location-anonymisation) and the age of the person stopped has been adjusted to a corresponding age group (e.g.
+18-24).
+In this research, 2.6 million stop and search incidents have been compiled for the observation period from January 2019 to December 2023, for seven legislative grounds (three of which are listed in Table 5) for 11 types of Searched Objects.
+Upon the aggregation of the data into three listed categories, drug objects related stop and search incidents took up 64% over the observation period, as compared to 12% for acquisitive objects and 24% for expressive objects.
+The searched subjects’ ethnic information have been collected from “officer defined ethnic group”.
+Table 5: Legislative bases underlying the aggregation of stop and search interactions
+Category Searched Objects Legislation
+Acquisitive objects ‘Stolen goods’, ‘Article for use in theft’ ‘Police and Criminal Evidence Act 1984 (section 1)’, ‘Criminal Justice Act 1988 (section 139B)’, ‘Police and Criminal Evidence Act 1984 (section 6)’
+Expressive objects ‘Offensive weapons’, ‘Anything to threaten or harm anyone’, ‘Firearms’, ‘Fireworks’, ‘Evidence of offences under the Act’, ‘Crossbows’, ‘Game or poaching equipment’ ‘Police and Criminal Evidence Act 1984 (section 1)’, ’Criminal Justice Act 1988 (section 139B)’, ‘Police and Criminal Evidence Act 1984 (section 6)’, ‘Criminal Justice and Public Order Act 1994 (section 60)’, ‘Firearms Act 1968 (section 47)’
+Drug objects ‘Controlled drugs’, ‘Psychoactive substances’ ‘Misuse of Drugs Act 1971 (section 23)’, ‘Psychoactive Substances Act 2016 (s36(2))’